OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -2 and 16, respectively.

OpenStudy (anonymous):

i think it's an = 1 • (-2)n + 1

OpenStudy (anonymous):

for a geometric series, we have \[a_n=a_0\cdot r^{n-1}\\ a_3=-2=a_0r\\ a_5=16=a_0 r^4 \] find out "r" first, then get the "\(a_0\)" and find the required term

OpenStudy (anonymous):

is it an = 1 • 2n

OpenStudy (anonymous):

is what \(1\cdot 2n\) ?

OpenStudy (anonymous):

yes 1 x 2^n

OpenStudy (anonymous):

the equation that you first wrote does not belong to "arithmatic" nor "geometric" sequence.

OpenStudy (anonymous):

ahhhh im confused

OpenStudy (anonymous):

did you understand the above formulae?

OpenStudy (anonymous):

a little i tried it again and got an = 1 • (-2)^n - 1

OpenStudy (anonymous):

is this right @electrokid